CAT Chat Minutes
March 24, 2003
Information and Follow-up
Paul Zschack said he had a situation at 34ID last week with kinematic mounts under critical components (white beam mask) in FOE. There is a potential shifting of component position due to unequal vacuum pressures. The component should be labeled individually as a critical component and the unequal pressure issue addressed. Did anyone else have this problem?
Roger replied that this was an APS responsibility and hell look into the problem.
User Operations
Machine Status R. Klaffky: Roger presented the following stats as of 3/24/03
Delivered beam: 1075.96 Hours
Percentage of scheduled time (*) 97.37 %
Downtime during period: 29.06 Hours
Percentage of scheduled time SR current > 10 ma 98.37%
Average delivered current during period: 97.49 ma
Mean time between faults (MTBF): 38.43 Hours
Faults per day of delivered beam: 0.62
Beamline Faults: There were three reported machine faults over the last week:
| # 68 03/18 16:00 to 03/20 10:06 42.10 | RF4 Power loss [RF] | 0.29 Refill |
| # 69 03/20 10:24 To 03/21 12:26 26.04 | 2ID PSS trip [ES] | 0.61 Magnet Standardize, refill |
| # 70 03/21 13:02 To 03/22 04:07 15.08 | RF4 cav.Vac.Flt [RF] | 0.18 Refill |
| # 71 03/22 04:18 To 03/24 07:59 51.69 | Int Dump: End of Period | 0.00 |
Girder Displacement Schedule O. Singh: Om presented an updated schedule of girder displacements (follows):
|
Displaced Dates |
To be displaced schedule |
Cleaned Xbpms ID |
Comments |
1 |
No schedule |
|||
2,3,4 |
October, 2000 |
2,3 |
||
5 |
January, 2001 |
4 |
||
6,7,8 |
April, 2001 |
5,6,7 |
||
9,10 |
January, 2003 |
8,9 |
||
11,12,13 |
April, 2003 |
10,11,12 |
||
14 |
January, 2004 |
13,14 |
||
15,16,17 |
May, 2002 |
15,16 |
||
18,19,20 |
September, 2002 |
17,18,19 |
||
21 |
January, 2003 |
20,21 |
||
22,23 |
January, 2002 |
22 |
||
24 |
January, 2004 |
23 |
||
25,26,27 |
May, 2004 |
24,25,26 |
||
28,29,30 |
No schedule |
27,28,29,30 |
||
31 |
January, 2004 |
31 |
||
32 |
July, 2000 |
32 |
||
33 |
April, 2000 |
33 |
||
34,35 |
January, 1999 |
34 |
||
36 |
No schedule |
|||
37,38,39,40 |
No schedule |
RF systems - no BL |
Note 1 - Next Shutdown schedule - April/May 2003 sectors 11,12,13
Note 2 - No Girder displacement work during September, 2003 shutdown due to other Survey group commitments.
This schedule is also displayed on the APS website at url: http://www.aps.anl.gov/asd/diagnostics/xaxstf/distortSched.html
Shipments of Radioactive Material & Foreign Soil B. Glagola: Bruce explained that DOE has stopped all radioactive material shipments. If you are expecting a shipment, and it is in process, you should have no problem receiving it but will not be able to send back. If you have any questions, please contact Bruce at 2-9797.USDA is requiring special permit to allow foreign soil brought into ANL. Bruce is in the process of obtaining a permit for APS to allow foreign soil in/out of ANL. Below is Bruces presentation and explanation on this USDA procedure:

Page 1 of 5

Page 2 of 5

Page 3 of 5

Page 4 of 5

Page 5 of 5
If you have questions or concerns, contact Bruce Glagola at 2-9797.